OpenStudy (calculusxy):

Solving Systems of Equations by Substitution. 1. -3x + 3y = 4 -x + y = 3 2. -5x - 8y = 17 2x - 7y = -17

OpenStudy (calculusxy):

@jim_thompson5910 @Nnesha

OpenStudy (calculusxy):

@mathstudent55

OpenStudy (calculusxy):

This is my work for question 1. we could figure out the value of y by: y = 3 + x -3x + 3(3 + x) = 4 is that correct?

OpenStudy (calculusxy):

i need to solve for x, which i know, but i just wanted to know whether my set up was fine or not.

Nnesha (nnesha):

yep yep right :=)

OpenStudy (calculusxy):

so then we can simply it to: -3x + 9 + 3x = 4

Nnesha (nnesha):

yep

OpenStudy (calculusxy):

but if we do -3x + 3x = 0x and I am not even sure if there is any such thing as 0x.

Nnesha (nnesha):

yes so they cancel out

Nnesha (nnesha):

thats mean no solution both sides are not equal \[9 \cancel =4\]

OpenStudy (calculusxy):

oh okay :)

OpenStudy (calculusxy):

and for the next one...

Nnesha (nnesha):

yep

OpenStudy (calculusxy):

we can figure out the value of x... 2x - 7y = -17 2x = -17 + 7y 2x / 2 = (-17 + 7x) / 2 x = -8.5 + 3.5x

Nnesha (nnesha):

hmm you mean x = -8.5 +3.5 y right ?

OpenStudy (calculusxy):

yes sorry :)

Nnesha (nnesha):

its okay yep right

OpenStudy (calculusxy):

-5(-8.5 + 3.5y) - 8y = 17

OpenStudy (calculusxy):

42.5 - 17.5y - 8y = 17 42.5 - 25.5y = 17 -25.5 - 25.5y = 17 -25.5y = 42.5 y = 1.67 i am pretty sure i am wrong it's going on decimals.

Nnesha (nnesha):

well i don't understand the 3rd line

OpenStudy (calculusxy):

if i did 17 - 42.5 = -25.5

Nnesha (nnesha):

yep that's right you need to move 42.5 to the right side then you will get Perfect one

OpenStudy (calculusxy):

what do you mean by moving 42.5?

Nnesha (nnesha):

i meant you have to subtract 42.5 both sides :=)

OpenStudy (jacob902):

correct

OpenStudy (calculusxy):

so what i did was correct right?

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @calculusxy 42.5 - 17.5y - 8y = 17 42.5 - 25.5y = 17 -25.5 - 25.5y = 17 -25.5y = 42.5 y = 1.67 i am pretty sure i am wrong it's going on decimals. \(\color{blue}{\text{End of Quote}}\) well here 3rd line you need to subtract 42.5 both sides i don't why there is -25.5 -25.5y

OpenStudy (calculusxy):

oh yeah

OpenStudy (calculusxy):

my mistake :\

Nnesha (nnesha):

:=) doing gREAT

OpenStudy (calculusxy):

y = 1 right?

Nnesha (nnesha):

yep right!

OpenStudy (calculusxy):

and then we would do: -5x - 8(1) = 17 -5x - 8 = 17 -5x = 25 x = -5

Nnesha (nnesha):

gO_OD to go!

OpenStudy (calculusxy):

thanks :) I have another question. it's small and related to basic physics. can u help?

Nnesha (nnesha):

PHYSICS hmmmm not sure

OpenStudy (calculusxy):

can u just see it? if u can't do it then u don't have to.

Nnesha (nnesha):

ye sure!

OpenStudy (calculusxy):

okay. here it is...

OpenStudy (calculusxy):

If i'm driving in a car at 30kph and a friend is driving in a car with the same mass at 60kph, my friend's car's kinetic energy is bigger than mine by a factor of what? (that is, divide my friend's over mine.)

Nnesha (nnesha):

i'm not sure havn't done physics n a while sorry

OpenStudy (calculusxy):

okay thank you though :)

Nnesha (nnesha):

np good job!